a) You are provided with the following information relating to V ltd


Equity and liabilities


12% debentures (shs1000 at par) 16,000


10% preferences shares 6,250


Ordinary shares (Shs 10 par) 12,500


Retained earnings 28,125


Additional information


i. The debentures are currently selling at Shs 950 in the market


ii. Company paid a dividend of Shs 5.00 per ordinary share and they are expected to grow at a rate of 10% per annum.


iv. The corporation tax is 40%


Required


Effective Cost of debt (3 marks)


Cost of equity (3 marks)


Weighted Average cost of capital (4 marks)



1
Expert's answer
2022-03-07T17:30:09-0500

1) Effective Cost of Debt Solution:

"12(1-0.4)=7.2"


2)Cost of equity

"28125+12500+6250=46875"


3)"WACC=12\frac{16000}{16000+46875}(1-0.4)+10\frac{6250}{16000+46875}+10\frac{12500+28125}{16000+46875}=9.29"